OpenStudy (anonymous):

a 11.0 L sample of hydrogen gas has a pressure of 25.0atm. What volume would this gas occupy at 2.30 atm?

OpenStudy (kc_kennylau):

PV=nRT When P=25, V=11: 11*25=nRT nRT=275 When P=2.3, 2.3V=nRT 2.3V=275 V=275/2.3=119.57L

OpenStudy (kc_kennylau):

I was never taught the presentation, so follow the presentation of your book (I learnt it all by myself and my dad, so there were no books)

OpenStudy (anonymous):

yea i figured it out already, thanks a lot tho =]

OpenStudy (kc_kennylau):

no problem :) (PostScript: Hey, can you teach me the presentation plz? orz)